# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 3.2 Analysis of the micro model
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\leq(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}.$
@@
$\begin{split}\left|(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}\right|&\leq\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\int_{0}^{t}\int_{\Gamma_{\varepsilon}}|\theta_{\varepsilon,k}^{s}|\,\mathrm{d}x\,\mathrm{d}\tau\\ &\leq C\|f_{\varepsilon}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\left(\|\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{1}((0,t)\times\Omega^{s}_{\varepsilon})}\right)\\ &\leq\|f_{\varepsilon}\|^{2}_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}C_{\delta}\|1\|_{L^{2}((0,t)\times A_{\varepsilon,k}^{s})}^{2}+C\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+\delta\|\nabla\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}.\end{split}$ (3.6)
@@
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right]$
Isolating the terms that are connected to the temperature of the solid, we have
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|A_{\varepsilon,k}^{s}(\tau)|\,\mathrm{d}\tau\right].$ (3.7)
With the transformation $\Phi$ from Lemma 1 we can switch to the fixed domain $\Omega$. Note that $x\in\Omega\setminus\Phi(A_{\varepsilon,k}^{s}(\tau))$ iff $\theta_{\varepsilon,k}^{s}\circ\Phi(x,\tau)=0$ for a.a. $\tau\in(0,t)$. And since $\det{(\nabla\Phi)}(x)=({H-\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})})/{H}>0$, we conclude from Eq. 3.7, that also
$\|(\theta_{\varepsilon,k}^{s}\circ\Phi)(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\|\nabla(\theta_{\varepsilon,k}^{s}\circ\Phi)\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}$ $\leq C\left[\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right].$
By Gronwall’s inequality, we arrive at
$\|\theta_{\varepsilon,k}^{s}\circ\Phi\|_{L^{\infty}(S;L^{2}(\Omega))\cap L^{2}(S;H^{1}(\Omega))}\leq Ck\left(\int_{S}|\Phi(A_{\varepsilon,k}^{s}(\tau))|\,\mathrm{d}\tau\right)^{\frac{1}{2}}$
The domain $\Omega$ is independent of $\varepsilon$, therefore applying the results of [41, Chapter 2, Theorem 6.1] we obtain $\|\theta^{s}_{\varepsilon}\circ\Phi\|_{L^{\infty}(S\times\Omega)}\leq C$, where $C>0$ can be chosen independent of $\varepsilon$. By the properties of $\Phi$ this immediately leads to $\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}\leq C$ and by the properties of the trace also to $|\theta^{s}_{\varepsilon|\Gamma_{\varepsilon}}|=|(\theta^{s}_{\varepsilon}\circ\Phi)_{|\Sigma}|\leq C$.
For the estimate in the fluid domain, test with $(0,\theta^{f}_{\varepsilon,k})$, with similar computations as above we arrive at
$\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}+\alpha\|\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(\theta^{s}_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}$
The right-hand side can bounded by the same steps as in Eq. 3.6. Note, that the trace estimate from Lemma 2 introduces an additional $\nicefrac{{1}}{{\varepsilon}}$. Setting $k\geq\|\theta^{s}_{\varepsilon}\|_{L^{\infty}(S\times\Omega^{s}_{\varepsilon})}$, we have
@@
With the extension operator $\mathbb{E}_{\varepsilon}$ from Lemma 1, we transform to the non-rough thin layer. Define $\mathbb{E}_{\varepsilon}{A}^{f}_{\varepsilon,k}(\tau)=\{x\in\Omega_{\varepsilon}:\mathbb{E}_{\varepsilon}\theta^{f}_{e,k}(t,x)\neq 0\}$, clearly $|A^{f}_{\varepsilon,k}(\tau)|\leq|\mathbb{E}_{\varepsilon}A^{f}_{\varepsilon,k}(\tau)|$. Multiplying by $\varepsilon$, we obtain that on the non-rough thin layer, it holds
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega_{\varepsilon})}+\kappa^{f}\|\nabla\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega_{\varepsilon})}\leq C\left[\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right],$
@@
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\leq Ck\left(\int_{S}|\mathbb{E}_{\varepsilon}A_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right)^{\frac{1}{2}}.$
Applying again [41, Chapter 2, Theorem 6.1], we now have to be careful regarding the $\varepsilon$ dependency of the estimate. First we obtain
$\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C\left(1+C\beta_{\varepsilon}^{\frac{d+2}{2}}|S|^{\frac{1}{2}}|\Omega_{\varepsilon}|^{\frac{1}{2}}\right)\leq C\left(1+C\varepsilon^{\frac{1}{2}}\beta_{\varepsilon}^{\frac{d+2}{2}}\right),$
where $\beta_{\varepsilon}$ is the embedding constant from Lemma 3. By Lemma 3, $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{-\frac{1}{d+2}})$. Therefore $\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C$ for $C>0$ independent of $\varepsilon$. Lastly, since $\|\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}=\|\mathbb{E}_{\varepsilon}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\leq C$, we have the desired estimate. ∎
@@
Let the same assumptions be fulfilled as in Lemma 4 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 3.1, with $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$. If $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$, then there exists a $C>0$ independent of $\varepsilon$ such that
$\begin{split}\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})}\leq C\quad\text{ and }\quad\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime})}\leq C\sqrt{\varepsilon}.\end{split}$ (3.8)
@@
By testing Eq. 3.1 with arbitrary $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ one obtains for almost all $t\in S$
$\begin{split}\bigl{|}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}\bigr{|}\\ &\leq|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}|+\frac{1}{\varepsilon}|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}|+\alpha|(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}|+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\leq\kappa^{s}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{s}_{\varepsilon})}\|\nabla\varphi^{g}\|_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{\kappa^{f}}{\varepsilon}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(\Omega^{f}_{\varepsilon})}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}(\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}+\|\varphi^{s}\|_{L^{2}(\Gamma_{\varepsilon})})\\ &\hskip 20.0pt+C\|f_{\varepsilon}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{s}\|_{H^{1}(\Omega^{s}_{\varepsilon})}+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|.\end{split}$
Choosing $\varphi^{f}=0$, integrating over the time interval $S$, followed by Hölder’s inequality and applying (3.3) gives the desired estimate for $\partial_{t}\theta_{\varepsilon}^{s}$.
For estimating $\partial_{t}\theta^{f}_{\varepsilon}$ we set $\varphi^{s}=0$ then the only problematic part is the convection term (the trace over $\Gamma_{\varepsilon}$ can be estimated by Lemma 2). By Eq. 3.5 and $\operatorname{div}{v_{\varepsilon}}=0$
$\begin{split}\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}=-\frac{1}{\varepsilon}(v_{\varepsilon}\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}.\end{split}$
@@
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|v_{\varepsilon}\|_{L^{\infty}(S\times\partial\Omega^{f}_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\phi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}.\end{split}$
@@
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{C}{\varepsilon}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.\end{split}$
@@
The additional assumption $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$ is not a new requirement to the problem, as the solution of the Stokes problem will fulfill this estimate naturally.
@@
Let the Assumptions (A2) and (A5) be fulfilled. Then for all $\varepsilon>0$ and $\psi_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ there exists a unique solution $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 3.2. In addition, it holds
$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\sqrt{\varepsilon}\|\nabla u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq C\quad\text{ and }\quad\sqrt{\varepsilon}\|p_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq\frac{C}{\varepsilon}$ (3.9)
@@
Given the bounds of the viscosity (A2) and Assumption (A5), this is analog to the steps in [12, Section 3.2 and 3.4]. ∎
We introduce the operators $\mathbb{T}_{\varepsilon}:V_{\varepsilon}\to L^{2}(S\times\Omega)$, which maps a prescribed velocity field to the corresponding temperature field via Eq. 3.1, and the Stokes operator $\mathbb{S}_{\varepsilon}:L^{2}(S\times\Omega)\to V_{\varepsilon}$ of Eq. 3.2, which returns the velocity field for a given temperature distribution. We are looking for fixed points of the operator
$\mathbb{F}_{\varepsilon}:L^{2}(S\times\Omega)\to L^{2}(S\times\Omega),\quad\psi_{\varepsilon}\mapsto\mathbb{T}_{\varepsilon}\left(\mathbb{S}_{\varepsilon}(\psi_{\varepsilon})\right).$ (3.10)

### v3

@@ 3.2 Well-posedness and a prior estimates
$\begin{split}\frac{1}{2}\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}\\ &+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\leq(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}.\end{split}$ (12)
@@
$\begin{split}\left|(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}\right|&\leq\|f_{\varepsilon}^{f}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\int_{0}^{t}\int_{\Gamma_{\varepsilon}}|\theta_{\varepsilon,k}^{f}|\,\mathrm{d}x\,\mathrm{d}\tau\\ &\leq\frac{C}{\varepsilon}\|f_{\varepsilon}^{f}\|_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\left(\|\theta_{\varepsilon,k}^{f}\|_{L^{1}((0,t)\times\Omega^{f}_{\varepsilon})}+\|\nabla\theta_{\varepsilon,k}^{f}\|_{L^{1}((0,t)\times\Omega^{f}_{\varepsilon})}\right)\\ &\leq\frac{C_{\delta}}{\varepsilon}\|f_{\varepsilon}^{f}\|^{2}_{L^{\infty}((0,t)\times\Gamma_{\varepsilon})}\|1\|_{L^{2}((0,t)\times A_{\varepsilon,k}^{f})}^{2}+\frac{C}{\varepsilon}\|\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\frac{\delta}{\varepsilon}\|\nabla\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}.\end{split}$ (13)
@@
With the extension operator $\mathbb{E}_{\varepsilon}^{f}$ from Lemma 1, we transform to the non-rough thin layer. Define $B^{f}_{\varepsilon,k}(\tau)\coloneqq\{x\in\Omega_{\varepsilon}:\mathbb{E}_{\varepsilon}^{f}\theta^{f}_{e,k}(\tau,x)\neq 0\}$, then clearly $|A^{f}_{\varepsilon,k}(\tau)|\leq|B^{f}_{\varepsilon,k}(\tau)|$. Multiplying by $\varepsilon$, we obtain that on the non-rough thin layer, it holds
$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega_{\varepsilon})}+\|\nabla\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega_{\varepsilon})}\leq C\left[\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{2}((0,t)\times\Omega_{\varepsilon})}^{2}+k^{2}\int_{0}^{t}|B_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right],$
@@
$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S;L^{2}(\Omega_{\varepsilon}))\cap L^{2}(S;H^{1}(\Omega_{\varepsilon}))}\leq Ck\left(\int_{S}|B_{\varepsilon,k}^{f}(\tau)|\,\mathrm{d}\tau\right)^{\frac{1}{2}}.$
Applying [37, Chapter 2, Theorem 6.1], we now have to be careful regarding the $\varepsilon$ dependency of the estimate. First we obtain
$\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon,k}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C\left(1+C\beta_{\varepsilon}^{\frac{d+2}{2}}|S|^{\frac{1}{2}}|\Omega_{\varepsilon}|^{\frac{1}{2}}\right)\leq C\left(1+C\varepsilon^{\frac{1}{2}}\beta_{\varepsilon}^{\frac{d+2}{2}}\right),$
where $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{-\frac{1}{d+2}})$ is the embedding constant from Lemma 3. Therefore $\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega_{\varepsilon})}\leq C$ independent of $\varepsilon$. Consequently $\|\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}=\|\mathbb{E}_{\varepsilon}^{f}\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Omega^{f}_{\varepsilon})}\leq C$, and $\|\theta_{\varepsilon}^{f}\|_{L^{\infty}(S\times\Gamma_{\varepsilon})}\leq C$.
Plugging this result back into Equation 12 and choosing $k$ large enough, we obtain
$\|\theta_{\varepsilon,k}^{s}(t)\|_{L^{2}(\Omega^{s}_{\varepsilon})}=0\quad\text{ for almost all }t\in S,$
resulting in $\theta_{\varepsilon,k}^{s}=0$ in $L^{2}(S\times\Omega^{s}_{\varepsilon})$ and consequently $\theta^{s}_{\varepsilon}\in L^{\infty}(S\times\Omega^{s}_{\varepsilon})$.
As mentioned at the beginning, the estimate for $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}=0$ follows by the same steps only with the roles of $\theta^{s}_{\varepsilon}$ and $\theta^{f}_{\varepsilon}$ switched, the other case is also slightly simpler since the estimates in the solid domain do not depend on $\varepsilon$. Adding the solutions for both cases yields the solution for the general case $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}\neq 0$ and consequently the general $L^{\infty}$ estimate. ∎
@@
Let the same assumptions be fulfilled as in Lemma 5 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 7, with $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$. If $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$, then there exists a $C>0$ independent of $\varepsilon$ such that
$\begin{split}\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})}\leq C\quad\text{ and }\quad\|\partial_{t}\theta_{\varepsilon}\|_{L^{2}(S;H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime})}\leq C\sqrt{\varepsilon}.\end{split}$ (14)
@@
By testing Eq. 7 with arbitrary $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ one obtains for almost all $t\in S$
$\begin{split}&\bigl|\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}\bigr|\\ &\leq|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}|+\frac{1}{\varepsilon}|(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}|+\alpha|(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}|+|(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}}|+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\leq\kappa^{s}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{s}_{\varepsilon})}\|\nabla\varphi^{g}\|_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{\kappa^{f}}{\varepsilon}\|\nabla\theta_{\varepsilon}\|_{L^{2}(\Omega^{f}_{\varepsilon})}\|\nabla\varphi^{f}\|_{L^{2}(\Omega^{f}_{\varepsilon})}+|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}|\\ &\hskip 10.0pt+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}(\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}+\|\varphi^{s}\|_{L^{2}(\Gamma_{\varepsilon})})+C\|f_{\varepsilon}^{s}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{s}\|_{H^{1}(\Omega^{s}_{\varepsilon})}+\|f_{\varepsilon}^{f}\|_{L^{2}(\Gamma_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(\Gamma_{\varepsilon})}.\end{split}$
Choosing $\varphi^{f}=0$, integrating over the time interval $S$, followed by Hölder’s inequality and applying (9) gives the desired estimate for $\partial_{t}\theta_{\varepsilon}^{s}$.
For estimating $\partial_{t}\theta^{f}_{\varepsilon}$ we set $\varphi^{s}=0$ then the only problematic part is the convection term (the trace over $\Gamma_{\varepsilon}$ can be estimated by Lemma 2). By Eq. 11 and $\operatorname{div}{v_{\varepsilon}}=0$
$\begin{split}\frac{1}{\varepsilon}(v_{\varepsilon}\cdot\nabla\theta_{\varepsilon},\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}=-\frac{1}{\varepsilon}(v_{\varepsilon}\theta_{\varepsilon},\nabla\varphi^{f})_{S\times\Omega^{f}_{\varepsilon}}+\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}.\end{split}$
@@
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{1}{\varepsilon}\|v_{\varepsilon}\|_{L^{\infty}(S\times\partial\Omega^{f}_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}.\end{split}$
@@
$\begin{split}\left|\frac{1}{\varepsilon}(v_{\varepsilon}\cdot n\theta_{\varepsilon},\varphi^{f})_{S\times\partial\Omega^{f}_{\varepsilon}}\right|\leq\frac{C}{\varepsilon}\|\theta_{\varepsilon}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\|\varphi^{f}\|_{L^{2}(S\times\Sigma^{\text{out}}_{\varepsilon})}\leq\frac{C}{\sqrt{\varepsilon}}\|\varphi^{f}\|_{L^{2}(S;H^{1}(\Omega^{f}_{\varepsilon}))}.\end{split}$
@@
The additional assumption $\|v_{\varepsilon}\|_{L^{2}(S\times\Omega^{f}_{\varepsilon})}\leq C\sqrt{\varepsilon}$ is not a new requirement to the problem, as the solution of the Stokes problem satisfies this estimate naturally.
@@
Let the Assumptions (A2) and (A5) be fulfilled. Then for all $\varepsilon>0$ and $\psi_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ there exists a unique solution $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 8. In addition, it holds
$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\sqrt{\varepsilon}\|\nabla u_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq C\quad\text{ and }\quad\sqrt{\varepsilon}\|p_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}\leq\frac{C}{\varepsilon}$ (15)
@@
Existence and uniqueness follow from [42, Theorem 5.1], where the proof for constant viscosity is stated, which is transferable to our setup given the bounds assumed in (A2). Furthermore, by Assumption (A5), and the extension of the boundary function $U_{e}$, the velocity estimates follow by testing with $u_{\varepsilon}-U_{\varepsilon}$ and the Poincaré inequality, as carried out in [25, Section 3.2]. The pressure estimate follows by combining the estimate of the velocity gradient with the results in [35, Lemma 4.7] for the Bogovskii operator in thin porous layers (which can be directly transferred to our setup). See also [34] for further studies of the Bogovskii operator in thin non-porous layers. ∎
Next, we introduce the operators $\mathbb{T}_{\varepsilon}:V_{\varepsilon}\to L^{2}(S\times\Omega)$, which maps a prescribed velocity field to the corresponding temperature field via Eq. 7, and the Stokes operator $\mathbb{S}_{\varepsilon}:L^{2}(S\times\Omega)\to V_{\varepsilon}$ of Eq. 8, which returns the velocity field for a given temperature distribution. We are looking for fixed points of the operator
$\mathbb{F}_{\varepsilon}:L^{2}(S\times\Omega)\to L^{2}(S\times\Omega),\quad\psi_{\varepsilon}\mapsto\mathbb{T}_{\varepsilon}\left(\mathbb{S}_{\varepsilon}(\psi_{\varepsilon})\right).$ (16)